thealmightykumar:

The mean of a normally distributed dataset is 12, and the standard deviation is 2. ____% of the data points lies between 8 and 16.

Vocaloid:

The mean is 12 and the standard deviation is 2 Notice: mean + 2(standard deviation) = 12 + 2(2) = 16 Mean - 2(standard deviation) = 12 - 2(2) = 8 So the boundaries 8 and 16 represent +/- 2 standard deviations from the mean Looking at the normal distribution curve, what % of values are within 2 standard deviations of the mean? See attached. https://i0.wp.com/statisticsbyjim.com/wp-content/uploads/2021/08/empirical_rule_graph2.png?fit=572%2C384&ssl=1

Vocaloid:

Review the empirical rule By definition 95% of the data lies within 2 standard deviations of the mean.